OpenStudy (anonymous):

subract \[(7+t-5t ^{2}+2t ^{8})-(1+2t-4t ^{2}+5t ^{8})\]

OpenStudy (anonymous):

7-1+t-2t-5t^2+4t^2+2t^8-5t^8

OpenStudy (anonymous):

can you do the rest by yourself?

OpenStudy (anonymous):

i dont even know how to do these

OpenStudy (anonymous):

i missed this lesson and i never caught up

OpenStudy (anonymous):

I group like terms and apply "-" to the 2nd ()

OpenStudy (anonymous):

what do you mean

OpenStudy (anonymous):

I took 7 and subtract 1, then took t and subtract 2t and so on

OpenStudy (anonymous):

so what the next step that i have to do

OpenStudy (anonymous):

just have to do like terms and that's it. 7-1=6 t-2t=-t and do the last two like terms

OpenStudy (anonymous):

5t^2-4t^2=it^2 and 2t^8-5t^8=-3t^7

OpenStudy (anonymous):

so how would i write out the solving answer

OpenStudy (anonymous):

almost good -5t^2+4t^2=-t^2 2t^8-5t^8=-3t^8

OpenStudy (anonymous):

6-t-t^2-3t^8

OpenStudy (anonymous):

that would be how i write the answer thanks you for the help

OpenStudy (anonymous):

you are welcome. you can use mathway.com http://mathway.com/answer.aspx?p=calg?p=(7+t −5tSMB072+2tSMB078)−(1+2t−4tSMB072+5tSMB078)?p=2?p=?p=?p=?p=?p=0?p=?p=0?p=?p=?p=Simplify